what is the better buy - 4 quarts of ice cream for $6.67 or 4 L for $6.99?
Answer by Theo(13342)

1 liter is equal to 1.05669 quarts.

using this ratio, then 4 liters is equal to 1.05669 * 4 = 4.22676 quarts.

for 6.67, you can buy 4 quarts.

for 6.99, you can buy 4 liters.

4 liters is equivalent to 4.22676 quarts.

therefore, for 6.99 you can buy 4.22676 quarts.

the cost per quart for 4 liters at 6.99 is therefore 6.99 / 4.22676 = 1.65 rounded to 2 decimal places.

the cost per quart for 4 quarts at 6.67 is therefore 6.67 / 4 = 1.67 rounded to 2 decimal places.

the 4 liters of ice cream is the better buy because it's cheaper on a cost per quart basis.